# Defense Strategies for Autonomous Multi-agent Systems: Ensuring Safety and Resilience Under Exponentially Unbounded FDI Attacks

### v1

@@ IV Numerical Simulations
In this section, we validate the proposed SAAR control strategies through numerical simulations of a heterogeneous autonomous MAS under EU-FDI attacks on both CIL and OL. Fig. 2 illustrates the communication topology among the agents. As seen, the autonomous MAS consists of four followers (agent 1 to 4) and four leaders (agent 5 to 8), with the dynamics of the followers and the leaders modeled by the following equations:
$\dot{x}_{1}$ $=\begin{bmatrix}-2&1&0\\ 0&-3&1\\ 0.5&0&-1\end{bmatrix}x_{1}+\begin{bmatrix}1&0&0\\ 0&1&0\\ 0&0&1\end{bmatrix}u_{1},$
$\dot{x}_{2}$ $=\begin{bmatrix}-1&0&0.5\\ 0&-2&1\\ 0.5&0&-0.5\end{bmatrix}x_{2}+\begin{bmatrix}0.5&1&0\\ 1&0.5&0\\ 0&0&1\end{bmatrix}u_{2},$
$\dot{x}_{3}$ $=\begin{bmatrix}-1&1&0\\ 0&-3&1\\ 0&0.5&-1\end{bmatrix}x_{3}+\begin{bmatrix}1&0&0\\ 0&1&0\\ 0&0&1\end{bmatrix}u_{3},$
$\dot{x}_{4}$ $=\begin{bmatrix}-1&0.5&0\\ 0.5&-1.5&0.5\\ -0.5&0&-2\end{bmatrix}x_{4}+\begin{bmatrix}1&0&0\\ 0&1&0\\ 0&0&1\end{bmatrix}u_{4},$
$\dot{x}_{5,6,7,8}$ $=\begin{bmatrix}0&-2&1\\ 2&0&1\\ -1&-1&0\end{bmatrix}x_{5,6,7,8}.$
@@
We inject the following EU-FDI attack signals on the CIL and OL:
$\begin{array}[]{cc}\gamma^{a}_{1}=\begin{bmatrix}5.5e^{0.11t}\\ 1.5e^{0.23t}\\ -6.6e^{0.08t}\end{bmatrix},&\gamma^{ol}_{1}=\begin{bmatrix}-1.2e^{0.40t}\\ 1.5e^{0.61t}\\ 2.7e^{0.15t}\end{bmatrix},\\[15.00002pt] \gamma^{a}_{2}=\begin{bmatrix}2.3e^{0.15t}\\ -8.7e^{0.42t}\\ 14.5e^{0.14t}\end{bmatrix},&\gamma^{ol}_{2}=\begin{bmatrix}3.3e^{0.06t}\\ -2.2e^{0.27t}\\ -1.7e^{0.21t}\end{bmatrix},\\[15.00002pt] \gamma^{a}_{3}=\begin{bmatrix}7.6e^{0.35t}\\ -9.7e^{0.20t}\\ -17.2e^{0.06t}\end{bmatrix},&\gamma^{ol}_{3}=\begin{bmatrix}2.8e^{0.24t}\\ -5.0e^{0.04t}\\ -1.8e^{0.08t}\end{bmatrix},\\[15.00002pt] \gamma^{a}_{4}=\begin{bmatrix}-2.9e^{0.16t}\\ 5.2e^{0.15t}\\ -7.7e^{0.10t}\end{bmatrix},&\gamma^{ol}_{4}=\begin{bmatrix}-5.2e^{0.04t}\\ 2.4e^{0.13t}\\ -2.1e^{0.18t}\end{bmatrix}.\end{array}$
These attack signals, which grow exponentially over time, are crafted to evaluate the system’s ability to adapt and maintain resilience in the face of evolving adversarial conditions. Select $U_{1,2,3,4}=I_{2}$, and $Q_{1,2,3,4}=3I_{2}$. The controller gain matrices $K_{i}$ and $H_{i}$ are found by solving (14)-(16).
We choose $d_{s}=0.2$ for collision avoidance. The performance of the proposed SAAR control strategies, considering these safety constraints, is demonstrated through a set of comparative simulation results contrasted with non-SAAR standard control protocols [20]. The three sub-figures in Fig. 3 illustrate the initial three-dimensional (3D) positions of the agents and compare the system responses under standard control and SAAR control following the attacks. Both EU-FDI attacks on the OL and CIL are initiated at $t=3\,\operatorname{s}$. Fig. 3(a) depicts the initial positions of the agents, while Fig. 3(b) presents the agents’ 3D positions at $t=4.50\,\operatorname{s}$ under the standard control approach, where the containment objective is not achieved post-attack. In contrast, Fig. 3(c) displays the agents’ 3D positions at $t=4.46\,\operatorname{s}$ under the SAAR control approach, demonstrating that the UUB containment objective is preserved despite the attacks.
Figure 3: Leader-follower 3D motion evolution: (a) $0\,\operatorname{s}$ snapshot. (b) Snapshot after the attacks using the standard control. (c) Snapshot after the attacks using the SAAR control.
@@
Fig. 4(a) illustrates the evolution of the containment error $e_{c}$ under conventional control. The dimensional components of $e_{c}$ diverge, indicating a failure in achieving containment. In contrast, Fig. 4(b) depicts the system response under SAAR control following the injection of the attacks. While brief spikes in the error reflect the system’s transient response to the attacks, the SAAR controller quickly compensates, causing the errors to stabilize and enter a steady state. This behavior demonstrates the resilience of the SAAR controller in maintaining UUB containment performance by effectively mitigating the impact of EU-FDI attacks on both CIL and OL.
@@
Fig. 5 depicts the evolution of the distances between followers, denoted as $d_{12},d_{13},d_{14},d_{23},d_{24},$ and $d_{34}$, over time. Following the attacks, a brief spike in the distances is observed as the system experiences a transient response. However, the SAAR controller quickly compensates, and the distances stabilize shortly after, ensuring collision avoidance among followers. Fig. 4(b) and Fig. 5 show that the system successfully maintains safety and achieves the UUB containment under EU-FDI attacks on both OL and CIL, validating the effectiveness of the proposed SAAR defense strategies.
@@ V Conclusion
In this study, we have developed a SAAR control framework to address the challenges posed by EU-FDI attacks on autonomous MASs. We have introduced an attack-resilient OL and integrated safety constraints through QP to achieve collision avoidance, ensuring system safety even under attack. Additionally, a compensational signal has been designed to counteract the impact of EU-FDI attacks on the CIL. Rigorous Lyapunov stability analysis and 3D simulations have demonstrated the SAAR controller’s effectiveness in maintaining both safety and resilience in practical autonomous MAS environments, where leader agents span safe regions to guide heterogeneous follower agents while addressing collision avoidance.

### v2

@@ IV Numerical Simulations
We validate the proposed SAAR control strategies through numerical simulations of a heterogeneous autonomous MAS under EU-FDI attacks on both CIL and OL. Fig. 3 illustrates the communication topology among the agents. As seen, the autonomous MAS consists of four followers (agents 1 to 4) and four leaders (agents 5 to 8), with the following dynamics:
[…]
@@
We consider the following EU-FDI attack signals on the CIL and OL:
$\begin{array}[]{cc}\gamma^{a}_{1}=\begin{bmatrix}2.5e^{0.07t}\\ 1.5e^{0.04t}\\ -6.6e^{0.08t}\end{bmatrix},\gamma^{ol}_{1}=\begin{bmatrix}-1.2e^{0.10t}\\ 1.5e^{0.17t}\\ 2.7e^{0.15t}\end{bmatrix},\par\gamma^{a}_{2}=\begin{bmatrix}2.3e^{0.05t}\\ -4.7e^{0.05t}\\ 11.5e^{0.04t}\end{bmatrix},\gamma^{ol}_{2}=\begin{bmatrix}3.3e^{0.06t}\\ -2.2e^{0.15t}\\ -1.7e^{0.12t}\end{bmatrix},\\[15.00002pt] \gamma^{a}_{3}=\begin{bmatrix}3.6e^{0.10t}\\ -4.7e^{0.09t}\\ -10.2e^{0.06t}\end{bmatrix},\gamma^{ol}_{3}=\begin{bmatrix}2.8e^{0.14t}\\ -5.0e^{0.04t}\\ -1.8e^{0.08t}\end{bmatrix},\par\gamma^{a}_{4}=\begin{bmatrix}-2.9e^{0.09t}\\ 5.2e^{0.06t}\\ -7.7e^{0.07t}\end{bmatrix},\gamma^{ol}_{4}=\begin{bmatrix}-5.2e^{0.04t}\\ 2.4e^{0.13t}\\ -2.1e^{0.14t}\end{bmatrix}.\end{array}$
Select $U_{1,2,3,4}=I_{2}$, and $Q_{1,2,3,4}=3I_{2}$. The controller gain matrices $K_{i}$ and $H_{i}$ are found by solving (13)-(15).
We choose $d_{s}=0.3$ for collision avoidance. The performance of the proposed SAAR control strategies, considering these safety constraints, is demonstrated through a set of comparative simulation results contrasted with non-SAAR standard control protocols [19]. The three sub-figures in Fig. 4 illustrate the initial 3D positions of the agents and compare the system responses under standard control and SAAR control following the attacks. Both EU-FDI attacks on the OL and CIL are initiated at $t=3\,\operatorname{s}$. Fig. 4(a) depicts the initial positions of the agents, while Fig. 4(b) presents the agents’ 3D positions at $t=15.08\,\operatorname{s}$ under the standard control approach, where the containment objective is not achieved after attack initiation. In contrast, Fig. 4(c) displays the agents’ 3D positions at $t=15.08\,\operatorname{s}$ under the SAAR control approach, demonstrating that the UUB containment objective is preserved despite the attacks.
Figure 4: Leader-follower 3D motion evolution: (a) $0\,\operatorname{s}$ snapshot. (b) $15.08\,\operatorname{s}$ snapshot (after attack initiation) using the standard control. (c) $15.08\,\operatorname{s}$ snapshot (after attack initiation) using the SAAR control.
@@
Fig. 5(a) illustrates the evolution of the containment error $e_{c}$ under conventional control. The dimensional components of $e_{c}$ diverge, indicating a failure in achieving containment. In contrast, Fig. 5(b) depicts the system response under SAAR control under attack injection. The SAAR controller quickly compensates after the attacks are initiated, causing the errors to stabilize and enter a steady state. This behavior demonstrates the resilience of the SAAR controller in maintaining UUB containment performance by effectively mitigating the impact of EU-FDI attacks on both CIL and OL.
@@
Fig. 6 depicts the evolution of the distances between followers, denoted as $d_{12},d_{13},d_{14},d_{23},d_{24},$ and $d_{34}$, over time, which shows the show safe distances among followers are maintained. Fig. 5(b) and Fig. 6 show that the system successfully maintains safety and achieves the UUB containment under EU-FDI attacks on both OL and CIL, validating the effectiveness of the proposed SAAR defense strategies.
@@ V Conclusion
This letter has formulated and solved the SAAR control problem for heterogeneous MASs to mitigate EU-FDI attacks on both OL and CIL while ensuring collision-free behaviors. An attack-resilient OL and a novel compensational signal on the CIL have been designed to address the EU-FDI attacks. A SAAR controller has then been developed by formulating and solving a QP problem integrating CBF certified collision-free safety constraints. Rigorous Lyapunov stability analysis and 3D simulations have demonstrated the SAAR controller’s effectiveness in maintaining both safety and resilience in practical autonomous MAS environments, where leader agents span safe regions to guide heterogeneous follower agents while ensuring collision avoidance.